OpenStudy (anonymous):

Challenge problem/discussion for you guys: Let say number \(e\) was never calculated... How can we find the value of \(a\) where \(f(x) = a^x = f'(x)\)?

OpenStudy (perl):

we are given that f(x) = f ' (x) now by definition f ' (x) = lim [f(x + h) - f(x)] /h = lim [a ^ ( x + h) - a^x ] / h = lim [a^x * a^h - a^x] / h = lim [a^x ( a^h -1 ) / h ] = a^x * [lim a^h -1 ] /h so we have so far a^x = a^x lim [ a^h -1 ] / h

OpenStudy (perl):

usually e is computed with an infinite series

OpenStudy (zzr0ck3r):

a=0

OpenStudy (anonymous):

Ok, is that only possible value of a?

OpenStudy (zzr0ck3r):

well of course e, and actually ae^x do we know that e^x is the inverse of ln?

OpenStudy (anonymous):

e? ln? What are you talking about? ;)

OpenStudy (zzr0ck3r):

\(\frac{dy}{dx}=y\implies \int \frac{1}{y}dy=\int dx\implies y\ne 0\)

OpenStudy (zzr0ck3r):

ahh ok will then that don't work:)

OpenStudy (perl):

micah , one moment

OpenStudy (anonymous):

haha

OpenStudy (perl):

ok so you agree with the derivative of f(x) = a^x

OpenStudy (zzr0ck3r):

I think you will need series:(

OpenStudy (anonymous):

Yeah, I agree?

OpenStudy (perl):

do you agree that for f(x) = a^x f ' (x) = lim [f(x + h) - f(x)] /h = lim [(a ^ ( x + h) - a^x ) / h] = lim [(a^x * a^h - a^x) / h ] = lim [a^x ( a^h -1 ) / h ] = a^x * lim [( a^h -1 ) / h ]

OpenStudy (zzr0ck3r):

I do

OpenStudy (anonymous):

Yeah @perl

OpenStudy (perl):

lets define a function m(a) = lim ((a^h - 1) /h) , as h->0 m(2) <1 , by evaluation m(3)>1 so by intermediate value theorem, and since m is a continuous function there exists a value , lets call it e, such that m(e) = 1 this simplifies the expression

OpenStudy (perl):

e is between 2 and 3

OpenStudy (perl):

now f ' (x) = m(a) * a^x , and if a = e , we have f ' (x) = m(e) * a^x = 1*a^x

OpenStudy (perl):

' by evaluation ' i mean numerical evaluation

OpenStudy (perl):

so yes, i actually was write before. we need to find a value 'a' such that lim (a^h -1 / h) = 1 , as h->0

OpenStudy (anonymous):

Yep, we know that a is in between 2 and 3, but I am sure we can do better than that. But I am stuck on \(\lim_{h\to0}\dfrac{a^h-1}{h}=1\) lol...

OpenStudy (perl):

you can use intermediate value theorem

OpenStudy (perl):

since that limit is continuous , lets call that limit m(a) , since it changes with different 'a'. there is some decimal 'a' such that we know that m(a) is continuous and m(2) < 1 < m(3) so there must be come decimal between 2 and 3 such that m(2) < 1= m(e) < m (3)

OpenStudy (anonymous):

Well, yeah, but it looks like it takes numerous of numerical evaluation for us to find good approximate of a. plus how did you evaluate m(x)? like you did with m(2) and m(3)

OpenStudy (perl):

you can make a table

OpenStudy (perl):

also i should add, a = 0 is undefined for m(a). but m(a) is continuous on [2,3]

OpenStudy (anonymous):

I am sorta new to calculus and I was just wondered how e came in. *groan* guess we would have to do that stuffs to calculate e, like how most of mathematicians in past calculated well known number such pi or gold ratio lol... Thanks, I am satisfied.

OpenStudy (anonymous):

Gonna left this question open, in case someone have different approaches. That's ok @perl ?

OpenStudy (perl):

ill make a table with a calculator, ok ?

OpenStudy (perl):

sure

OpenStudy (anonymous):

ok cool

OpenStudy (perl):

my draw button seems to be failing, so ill be back later when i get a chance

OpenStudy (perl):

basically you do h = .01 (2^.01 -1 ) / .01= h = .001 (2^.001 -1 ) / .001=

OpenStudy (perl):

that will converge to m(2), otherwise known as ln 2

OpenStudy (anonymous):

ah, and I takes that we should do evaluate on m(2.5) and see whether it is smaller than 1 or not, then evaluation m(2.75)? and so on, then we will approaches to e, right?

OpenStudy (perl):

correct, 2.75 is too big though

OpenStudy (perl):

so you will have to go back to 2.25

OpenStudy (anonymous):

m(2.5) should be smaller than 1, so we should then go next to m(2.625)

OpenStudy (anonymous):

basically just go half and half all the way.

OpenStudy (perl):

m(2) < 1 m(3) >1 then by bisection method, we choose the halfway point m(2.5) <1 then m(2.75 ) > 1 oops, too big then halfway between 2.5 and 2.75

OpenStudy (perl):

right

OpenStudy (anonymous):

yeah

OpenStudy (perl):

ok i see what youre doing, you want to actually determine e with some degree of accuracy.

geerky42 (geerky42):

Another approach if you are still interested @micahwood50: Since \(f(x) = f'(x)\), it is true that \(f'(x) = f''(x)\) and \(f''(x) = f'''(x)\) and so on... So we can see that \(f(x) = f^{(n)}(x)\quad\forall n\in\mathbb{N}\), right? Ok, we can expand \(a^x\) into infinite polynomial using Maclaurin Series. \[f(x) = \sum_{n=0}^\infty f^{(n)}(0)\dfrac{x^n}{n!}\] So... \[a^x = \sum_{n=0}^\infty a^0\dfrac{x^n}{n!} = \sum_{n=0}^\infty \dfrac{x^n}{n!}\] Note that \(\displaystyle \dfrac{\mathrm d}{\mathrm d x}\sum_{n=0}^\infty \dfrac{x^n}{n!} = \sum_{n=0}^\infty \dfrac{x^n}{n!}\) So finally we can calculate \(a\) by letting x=1: \[f(1) = a = \sum_{n=0}^\infty \dfrac{1}{n!}\] Add up to, like, first 7 terms, and you should have good approximate value of a.

OpenStudy (anonymous):

This limit actually is another definition for \(e\).

OpenStudy (anonymous):

And it just happens to be equivalent to other definitions of \(e\).

OpenStudy (anonymous):

But hypothetically, if it was a different value, then we would have defined another constant just for the sake of this limit, since this limit allows us to differentiate exponential functions.